OpenStudy (anonymous):

Need help on this problem. @jim_thompson5910

jimthompson5910 (jim_thompson5910):

What's the center of this circle?

OpenStudy (anonymous):

I'm not sure for this one.

jimthompson5910 (jim_thompson5910):

what are some points on this circle?

OpenStudy (anonymous):

-1, 2, 1,1 , 2,1

jimthompson5910 (jim_thompson5910):

(1,7) is another point

jimthompson5910 (jim_thompson5910):

(1,1) and (1,7) form a diameter what's the midpoint of this segment

OpenStudy (anonymous):

1,4

jimthompson5910 (jim_thompson5910):

that's the center of the circle

jimthompson5910 (jim_thompson5910):

What's the radius?

OpenStudy (anonymous):

I thought it was 1,4 because that's the middle - guess not :/

jimthompson5910 (jim_thompson5910):

what's the distance from the center (1,4) to any point on the circle?

OpenStudy (anonymous):

1, 4 to 1, 7 is 3

jimthompson5910 (jim_thompson5910):

center: (1,4) Radius: 3

jimthompson5910 (jim_thompson5910):

use those two facts to get the equation

OpenStudy (anonymous):

I plug in the center into the x and y's?

jimthompson5910 (jim_thompson5910):

the general circle equation is \[\Large (x-h)^2 + (y-k)^2 = r^2\]

jimthompson5910 (jim_thompson5910):

Center: (1,4) (h,k) = (1,4) h = 1 k = 4 plug those into \[\Large (x-h)^2 + (y-k)^2 = r^2\]

jimthompson5910 (jim_thompson5910):

also, plug r = 3 into \[\Large (x-h)^2 + (y-k)^2 = r^2\]

OpenStudy (anonymous):

it would be C !

jimthompson5910 (jim_thompson5910):

good